1. An object was launched at 300 angle and travel 55.75m. Assume That the object was launched on a level ground.


a. What was the take-off speed of the object?


1
Expert's answer
2020-12-14T07:17:46-0500

Find the horizontal component of velocity:


"v_x=v\text{ cos}30°."


55.75 m is the horizontal range. The time it takes to travel this distance at "v_x" is


"t=\frac{R}{v_x}=\frac{R}{v\text{ cos}30°}"

On the other hand, this time is equal to the time required for the body to reach the highest point and then fall to the ground:


"t=2\cdot\frac{v_y}{g}=\frac{2v\text{ sin}30°}{g}."

So, we have two equal time variables expressed differently.


"\frac{2v\text{ sin}30°}{g}=\frac{R}{v\text{ cos}30°},\\\space\\
v=\sqrt{\frac{gR}{\text{sin}(2\cdot30°)}}=25.12\text{ m/s}."